The product of two numbers is 28. If the second number is one more than twice the first number, what are the two sets of numbers?
Answer by mananth(16949)

let the numbers be x & y
xy=28
x=28/y
y-2x=1
subxtitute x
y-2(28/y)=1
y^2-56=y
y^2-y-56=0
y^2-8y+7y-56=0
y(y-8)+7(y-8)=0
(y-8)(y+7)=0
y=8 OR -7
xy = 28
x=-7
so y =-4
(-7,-4)
y=8
then x= 28/8
x=3.5
(3.5,8)